40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog, Page 7 of 9  Not logged in ELOG logo
ID Date Author Type Category Subject
  110   Sat Apr 13 21:06:02 2013 KojiOpticsGeneralOMC Bottom-side: cavity glued

[Jeff, Zach, Lisa, Koji]

Gluing of the cavity mirrors went very well!!!

Preparation

- Checked if the cavity is still resonating. => Yes.

- Checked the FSR: 264.251MHz => 1.1345m
  2.5mm too long => Move each micrometer by 0.625mm backward

- FSR&TMS (I)
  Aligned the cavity again and checked the FSR: 264.8485MHz => 1.13194m
  TMS(V): 58.0875MHz => gamma_V = 0.219324
  TMS(H): 58.1413MHz => gamma_H = 0.219526
  the 9th modes of the carrier is 9.7-10.4 line width (LW) away from the carrier resonance
  the 13th modes of the lower f2 sideband are 9.2-10.2 LW away
  the 19th modes of the upper f2 sideband are 0.3-1.8 LW away
  We found that this coincidence of the resonance can be corrected by shortening the cavity round-trip by 0.5mm

- Spot positions (I)
 
The spots on the curved mirrors were ~1mm too much inside (FM side). In order to translate the cavity axis,
  MM2 and MM4 were pushed by θ
  θ/2.575 = 1mm ==> θ = 2.6 mrad
  The separation of the micrometers are ~20mm
  d/20mm = 2.6mrad ==> d = 52um

  1div of the micrometer corresponds to 10um => 5div = 50um

- Move the micrometers and adjusted the input steering to recover the alignment.

- In any case we were confident to adjust the FSR/TMS/spot positions only with the micrometers

BS1/FM1/FM2 gluing

- Aligned the cavity

- Glued BS1/FM1/FM2 one by one while the cavity resonance was maintained.
  FM2 was slipping as the table is not leveled well and the fixture was not supporting the optic.

- FSR&TMS (II)
  FSR: 264.964875MHz => 1.13144m (Exactly 0.5mm shorter!)
  TMS(V): 58.0225MHz => gamma_V = 0.218982
  TMS(H): 58.1225MHz => gamma_H = 0.219359
  the 9th modes of the carrier is 10.3~11.7 LW away
  the 13th modes of the lower f2 sideband are 7.4~9.3 LW away
  the 19th modes of the upper f2 sideband are 1.5~4.4 LW away

- Spot positions (II)
  Looked OK.

CM2 gluing

- Glued CM2. The mirror was supported from the back with allen keys.

- FSR&TMS (III)
 
FSR: 264.9665625MHz => 1.13144m
  TMS(V): 58.1275MHz => gamma_V = 0.219377
  TMS(H): 58.0813MHz => gamma_H = 0.219202
  the 9th modes of the carrier is 10.2~10.9 LW away
  the 13th modes of the lower f2 sideband are 8.5~9.4 LW away
  the 19th modes of the upper f2 sideband are 1.4~2.7 LW away

- Spot positions (III)
  Looked slightly off at CM2. Pushed MM2 by 4um.

CM1 gluing

- Glued CM1.

- FSR&TMS (IV)
 
FSR: 264.964875MHz => 1.13144m
  TMS(V): 58.06625MHz => gamma_V = 0.219145
  TMS(H): 58.08625MHz => gamma_H = 0.219220
  the 9th modes of the carrier is 10.8~11.1 LW away
  the 13th modes of the lower f2 sideband are 8.2~8.6 LW away
  the 19th modes of the upper f2 sideband are 2.6~3.2 LW away

- Spot positions (final confirmation)
  Looked OK. 

Final measurement

- After everything was finished, more detailed measurement has been done.

- FSR&TMS (final)
 
FSR: 264.963MHz => 1.13145m
  TMS(V): 58.0177MHz => gamma_V = 0.218966
  TMS(H): 58.0857MHz => gamma_H = 0.219221
  the 9th modes of the carrier is 10.8~11.7 LW away
  the 13th modes of the lower f2 sideband are 7.3~8.6 LW away
  the 19th modes of the upper f2 sideband are 2.6~4.5 LW away

Final values for the micrometers

  • MM1: The one closest to the input mirror (CM1) 0.78mm
  • MM2: The other one on CM1 0.89
  • MM3: The one closest to the output mirror (CM2) 0.90
  • MM4: The other one on CM2 0.90
/------------\

0.90         0.78
\------------/
0.90         0.89

 

Attachment 1: Cav_scan_response_130412_Pitch.pdf
Cav_scan_response_130412_Pitch.pdf
Attachment 2: Cav_scan_response_130412_Yaw.pdf
Cav_scan_response_130412_Yaw.pdf
  109   Fri Apr 12 09:25:31 2013 KojiOpticsCharacterizationAlignment of the OMC (without glue)

[Zach Koji]

The first attempt not to touch the curved mirrors did not work. (Not surprising)
The eigenmode is not found on the mirror surface.

We decided to touch the micrometers and immediately found the resonance.
Then the cavity alignment was optimized by the input steering mirrors.

We got the cavity length L and f_TMS/f_FSR (say gamma, = gouy phase / (2 pi) ) as
    L=1.1347 m        (1.132m nominal)
    gamma_V = 0.219176    (0.21879 nominal)
    gamma_H = 0.219418    (0.21939 nominal)


This was already sufficiently good:
- the 9th modes of the carrier is away from the resonance 10-11 times
  of the line width (LW)
- the 13th modes of the lower f2 sideband are 9-10 LW away
But
- the 19th modes of the upper f2 sideband are 1-3 LW away
  This seems to be the most dangerous ones.
and
- The beam spots on the curved mirrors are too marginal

So we decided to shorten the cavity round-trip 2.7mm (= 0.675mm for each micrometer)
and also use the curved mirrors to move the eigenmode toward the center of the curved mirrors.

After the movement the new cavity length was 1.13209 m.
The spot positions on the curved mirrors are ~1mm too close to the outside of the cavity.
So we shortened the outer micrometers by 8um (0.8 div).
This made the spot positions perfect. We took the photos of the spots with a IR sensor card.

The measured cavity geometry is (no data electrically recorded)
    L=1.13207 m        (1.132m nominal, FSR 264.8175MHz)
    gamma_V = 0.218547    (0.21879 nominal, 57.8750MHz)
    gamma_H = 0.219066    (0.21939 nominal, 58.0125MHz)

- the 9th modes of the carrier is 11-13 LW away
- the 13th modes of the lower f2 sideband are 5-8 LW away
- the 19th modes of the upper f2 sideband are 4-8 LW away

The raw transmission is 94.4%. If we subtract the sidebands and
the junk light contribution, the estimated transmission is 97.6%.

Note:
Even if a mirror is touched (i.e. misaligned), we can recover the good alignment by pushing the mirror
onto the fixture. The fixture works pretty well!
 

  108   Thu Apr 11 15:10:22 2013 KojiGeneralGeneralOMC Progress

[Zach, Jeff, Koji]


- Jeff configured the bottom side template to have a nominal value
obtained from the solid works model. Note that the thickness of the
curved mirrors are 6mm in the model. He added 0.3mmx2 to the dimensions.

- Jeff located the template on the breadboard such that each side has
the same amount of hanging out.

- Micrometer values

  • The one closest to the input mirror (CM1) 0.07
  • The other one on CM1 0.24
  • The one closest to the output mirror (CM2) 0.17
  • The other one on CM2 0.30
/------------\

0.17         0.07
\------------/
0.30         0.24

- Now the template is ready to accept the OMC optics.

 


- Zach and Koji finished a series of measurements for the test OMC.

Modulation depth:

- We scanned the laser PZT and recorded the data.
CH1: Reflection DC
CH2: PDH Error
CH3: Transmission (Magnified)
CH4: Transmission

- We should be able to obtain the estimation of the modulation depth and the finesse from this measurement.

- Rough calculation of the modulation depth is 0.19

Transmission:

- Incident 16.3mW
- Transmission 15.1mW
- This gave us the raw transmission of 92.6%ish.
- The modulation depth of 0.19 corresponds to 1.8% of the incident power
- The carrier reflection is almost dominated by the mode mismatch. (Note: We did not have a good resolution for the refl beam)  =>3.2%

- In total:The incident useful carrier power was 15.4mW ==> Throughput 98%
- There is slight headroom to increase the transmission by cleaning the mirrors.

FSR/Finesse:

- As our AOM is not functioning now, phase modulation sidebands are injected with the BBEOM.
- In principle, we can't expect any signal at the transmission at around the FSR frequency.
- If we apply small locking offset, the split peaks appear at the FSR frequency. The frequency of the dip corresponds to the FSR.
- We probably can extract the finesse of the cavity from this measurement. Lisa is working on this.

HOM/Finesse:

- The same PM injection gives us the frequency of the HOMs.
- We found that our EOM can work until ~500MHz.
- We could characterize the cavity resonance structure more than a single FSR.

  107   Wed Apr 10 00:40:30 2013 ZachOpticsConfigurationfauxMC locked

[Koji, Zach]

Tonight, we locked the "fauxMC". We obtained a visibility of >99%.

Koji had aligned it roughly last night, but we wanted to have a couple steering mirrors in the path for this practice cavity (the periscope mirrors will serve this function in the real setup), so we marked the alignment with irises and installed two extra mirrors.

After obtaining flashes with the WinCam placed at the output coupler, we removed the WinCam and put a CCD camera at one of the curved mirror transmissions and used this to get a strong TEM00 flash. Then, we installed the REFL PD/CCD, swept the laser PZT and optimized the alignment by minimizing the REFL dips. Finally, we connected the RF electronics and locked the cavity with the LB box. We used whatever cables we had around to trim the RF phase, and then Koji made some nice SMA cables at the 40m.

One thing we noticed was that we don't have enough actuation range to keep the cavity locked for very long---even with the HV amp (100V). We are going to offload to the NPRO temperature using an SR560 or pomona box circuit. We may also make an enclosure for the cavity to protect it from the HEPA blasting.

Tomorrow, after we do the above things, we will practice measuring the transmission, length (FSR) and mode spectrum of the cavity before moving on to the real McCoy.

fauxMC_labeled.jpg

  106   Tue Apr 9 13:56:09 2013 KojiOpticsGeneralPZT assembly post gluing / pre baking pictures

 

 

  105   Mon Apr 8 23:42:33 2013 KojiOpticsConfigurationFake OMC roughly aligned

Mode matching:

 

Attachment 1: beam_profile.pdf
beam_profile.pdf
Attachment 2: mode_matching.pdf
mode_matching.pdf
  104   Mon Apr 8 21:11:14 2013 KojiOpticsGeneralPZT assembly gluing

[Jeff, Zach, Koji]

PZT assembly gluing

Glue gun -> to be returned to MIT
Fixtures x2
Al bases, spacers
spare screws
mirrors / prisms / PZTs
IPA bottle
clean tools x2
first contact kit
gloves (7.5)

  103   Mon Apr 8 20:56:52 2013 KojiOpticsConfigurationPZT & Curverd Mirror arrangement

Assembly #1:

Mounting Prism #16
PZT #26
Mirror C6

Assembly #2:

Mounting Prism #20
PZT #23
Mirror C5

Attachment 1: PZT_assembly.pdf
PZT_assembly.pdf PZT_assembly.pdf
  102   Mon Apr 8 11:49:18 2013 KojiMechanicsCharacterizationPZT actuator tested at LLO

Test result of the PZTs by Valera and Ryan

PZT  Length Angle
 #   [nm/V] [urad/um]
 11  14.5   17.6
 12  13.8   17.8
 13  11.2   25.0
 14  14.5    6.6
 15  12.5   10.6

 21  14.5    9.7
 22  13.8   28.8
 23  14.5    6.8  ==> Assembly #2
 24  18.5   51.7  ==> Used for prototyping
 25  17.1   13.8
 26  14.5    6.6  ==> Assembly #1
  101   Mon Apr 8 11:29:08 2013 KojiOpticsCharacterizationMirror/PZT Characterization links
  100   Mon Apr 8 11:11:37 2013 KojiOpticsCharacterizationMore Mirror T measurement

More Ts of the mirrors were measured.

A mirror specification:
Request: 8300+/-800 ppm
Data sheet: 7931ppm

C mirror specification:
Request: 50+/-10 ppm
Data sheet: 51.48ppm or 46.40ppm

 

Mirror | P_Incident P_Trans  P_Offset | T_trans
       | [mW]       [uW]     [uW]     | [ppm]
-------+------------------------------+---------
A1     | 10.28    82.9       -0.205   | 8.08e3
A2     | -----     -----     ------   | ------
A3     | 10.00    83.2       -0.205   | 8.34e3
A4     | 10.05    80.7       -0.205   | 8.05e3
A5     |  9.94    81.3       -0.205   | 8.20e3
A6     | 10.35    78.1       -0.205   | 7.57e3
A7     | 10.35    77.8       -0.205   | 7.54e3
A8     | 10.30    78.0       -0.205   | 7.60e3
A9     | 10.41    84.1       -0.205   | 8.10e3
A10    | 10.35    77.3       -0.205   | 7.49e3
A11    | 10.33    77.9       -0.205   | 7.56e3
A12    | 10.34    78.7       -0.205   | 7.63e3
A13    | 10.41    85.4       -0.205   | 8.22e3
A14    | 10.34    84.4       -0.205   | 8.18e3
-------+------------------------------+---------
C1     | 10.30     0.279     -0.225   | 48.9
C2     | -----     -----     ------   | ------
C3     | 10.37     0.240     -0.191   | 41.6
C4     | 10.35     0.278     -0.235   | 49.6
C5     | 10.40     0.138     -0.235   | 35.9 => PZT assembly #2
C6     | 10.34     0.137     -0.235   | 36.0 => PZT assembly #1
C7     | 10.37     0.143     -0.229   | 35.9
C8     | 10.41     0.224     -0.237   | 44.3
C9     | 10.36     0.338     -0.230   | 54.8
C10    | 10.39     0.368     -0.228   | 57.4
C11    | 10.38     0.379     -0.209   | 56.6
C12    | 10.28     0.228     -0.238   | 45.3
C13    | 10.36     0.178     -0.234   | 39.8
-------+------------------------------+---------

 

  99   Fri Apr 5 18:18:36 2013 ZachOpticsConfigurationAOM probably broken

Quote:

Then, I started to check the AOM path. I noticed that the 1st (or -1st) order beam is very weak.
The deflection efficiency is ~0.1%. Something is wrong.
I checked the driver. The driver's coupler output (1:10) show the amplitude ~1V. (good)
I check the main output by reducing the offset. When the coupler output is 100mV, the main output was 1V. (good)
So is the AOM itself broken???

As Koji noticed that the AOM efficiency was very low, I figured I would try looking at it with a fresh set of eyes. The end result is that I have to agree that the AOM appears to be broken.

First, I measured the input impedance of the AOM using the AG4395A with the impedance test kit (after calibrating). The plot is below. The spec sheet says the center frequency is 200 MHz, at which Zin should be ~50 ohms. It crosses 50 ohms somewhere near 235 MHz, which may be reasonable given that the LC circuit can be tuned by hand. However, it does surprise me that the impedance varies so much over the specified RF range of ±50 MHz. Maybe this is an indication that something is bad.

Isomet_1250C_zin_4_5_13.png

I removed the cover of the modulator (which I think Koji did, as well) and all the connections looked as I imagine they should---i.e., there was nothing obviously broken, physically.

I then tried my hand at realigning the AOM from scratch by removing and replacing it. I was not able to get better than 0.15%, which is roughly what Koji got.

So, perhaps our best course of action is to decide what we expect the Zin spectrum to look like, and whether that agrees with the above measurement.

  98   Fri Apr 5 14:39:26 2013 KojiMechanicsCharacterizationCalibration of the test PZTs after the heat cure

We attached fused silica windows on the test PZTs. http://nodus.ligo.caltech.edu:8080/OMC_Lab/93

The glued assemblies were brought to Bob's bake lab for the heat cure. There they are exposed to 94degC heat for two hours (excluding ramp up/down time).

After the heat cure, we made the visual inspection.
The photos are available here.

Pre-bake
Test PZT #1: 17.4nm/V
Test PZT #2: 17.2nm/V
Test PZT #3: 30.6nm/V

Post-bake
Test PZT #1: 27.2 nm/V
Test PZT #2: 26.9 nm/V
Test PZT #3: 21.4 nm/V

Measurement precision is ~+/-20%
Spec is 14nm/V

Attachment 1: shadow_sensor_calib_after_bake.pdf
shadow_sensor_calib_after_bake.pdf
Attachment 2: PZTresponse.pdf
PZTresponse.pdf
  97   Thu Apr 4 23:44:52 2013 KojiOpticsConfigurationBeam launched into fiber

We had to move our flipper mirror to share the beam between Peter's setup and ours as our flipper is at the place where the ISS PD array base is supposed to be!
There was no place to insert the flipper in the setup. We (Peter and Koji) decided to move the laser back for ~2".

This entirely changed the alignment of the setup. The fiber coupler was my reference of the alignment.
Once the beam is aligned, I check the coupling to the fiber. It was 50%.

I tweaked the lens and eventually the coupling is improved to 83%. (24.7mW incident, 20.4mW obtained.)

Then, I started to check the AOM path. I noticed that the 1st (or -1st) order beam is very weak.
The deflection efficiency is ~0.1%. Something is wrong.
I checked the driver. The driver's coupler output (1:10) show the amplitude ~1V. (good)
I check the main output by reducing the offset. When the coupler output is 100mV, the main output was 1V. (good)
So is the AOM itself broken???

  96   Thu Apr 4 01:43:06 2013 KojiOpticsCharacterizationMirror T measurement

[Zach, Koji]

The measurement setup for the transmission measurement has been made at the output of the fiber.

- First, we looked at the fiber output with a PBS. It wasn't P-pol so we rotated the ourput coupler.
  What we found was that it wasn't actually linearly polarized.
  So the input coupler was rotated to correct it. This terribly misaligned the input coupling.
  After some iteration of rotating and aligning the input/output couplers, we obtained reasonable
  extiction ratio like 10mW vs 100uW (100:1) with 11mW incidence. (Where is the rest 0.9mW!?)

- The P-pol (transmission) out of PBS goes into the mirror. Here we tested mirror A1.
  The mirror is mounted on the prism mount supported by a rotational stage for precise angle adjustment
  We limited the input power down to 5mW so that we can remove the attenuator on the power meter.
  The reading of the power meter was fluctuating, indeed depending on MY position.
  So we decided to turn off the lighting of the room. This made the reading very stable.

  The offset of the power meter was -0.58uW

  The transmitted power for the normal incidence was 39.7uW with the incident 4.84mW.
  [39.7-(-0.58)] / [4.84*1000-(-0.58)] *10^6 = 8320 ppm

  The transmitted power for the 4deg incidence was 38.0uW with the incident 4.87mW.
  [38.0-(-0.58)] / [4.87*1000-(-0.58)] *10^6 = 7980 ppm

 cf. The specification is 7931ppm

 

  95   Thu Apr 4 01:35:04 2013 KojiOpticsCharacterizationMode matching to the OMC cavity

The fiber output was matched with the lenses on a small bread board.
The detailed configuration is found in the following elog link.

http://nodus.ligo.caltech.edu:8080/OMC_Lab/105

  94   Thu Apr 4 00:35:42 2013 ZachOpticsConfigurationMMT installed on breadboard, periscope built

[Koji, Zach]

We installed the MMT that matches the fiber output to the OMC on a 6"x12" breadboard. We did this so that we can switch from the "fauxMC" (OMC mirrors arranged with standard mounts for practice locking) to the real OMC without having to rebuild the MMT.

The solution that Koji found was:

z = 0: front face of the fiber output coupler mount

z = 4.8 cm: f = 35mm lens

z = 21.6 cm: f = 125mm lens

This should place the waist at z ~ 0.8 m. Koji has the exact solution, so I will let him post that.

The lenses are on ±0.5" single-axis OptoSigma stages borrowed from the TCS lab. Unfortunately, the spacing between the two lenses is very close to a half-integer number of inches, so I had to fix one of them using dog clamps instead of the screw holes to preserve the full range.

Koji also built the periscope (which raises the beam height by +1.5") using a vertical breadboard and some secret Japanese mounts. Part of it can be seen in the upper left corner of the photo below---sorry for not getting a shot of it by itself.

2013-04-03_23.43.00.jpg

  93   Wed Apr 3 18:42:45 2013 KojiOpticsGeneralEP30-2 gluing test

EP30-2 gluing test

  92   Wed Apr 3 17:39:38 2013 KojiMechanicsCharacterizationCalibration of the test PZTs before the glue test

We want to make sure the responses of the PZT actuator does not change after the EP30-2 gluing.

A shadow sensor set up was quickly set-up at the fiber output. It turned out the ring PZTs are something really not-so-straightforward.
If the PZT was free or just was loosely attached on a plane by double-sided tape, the actuation response was quite low (30% of the spec).
After some struggle, I reached the conclusion that the PZT deformation is not pure longitudinal but some 3-dimensional, and you need to
use a "sandwitch" with two flat surfaces with some pressue.

I turned the setup for horizontal scans to the vertical one, and put the PZT between quarter-inch spacers.
Then two more spacers are placed on the stack so that the weight applies the vertical pressure on the PZT.
This is also use ful to adjust the height of the shadow.

P4033491.JPG

The calibration plot is attached. It gives us ~21k V/m.
Voltage swing of 150V results the output voltage change of ~50mV.  This is pretty close to what is expected from the spec (16nm/V).
The PZT#3 (which had the mirror glued on) showed significantly large response.

Test PZT #1: 17.4nm/V
Test PZT #2: 17.2nm/V
Test PZT #3: 30.6nm/V
UHV PZT #24: 17.6nm/V

These numbers will be checked after the heat cure of EP30-2

Attachment 2: shadow_sensor_calib.pdf
shadow_sensor_calib.pdf
  91   Mon Apr 1 18:17:01 2013 KojiOpticsGeneralMirror curvature center test

Locations of the curvature minimum on the OMC curved mirrors have been measured.

Motivation:

When a curved mirror is misaligned, the location of the curvature center is moved.
Particularly, our OMC mirror is going to be attached on the PZT and the mounting prism with the back surface of the mirror.
This means that a curved mirror has inherent misalignment if the curvature minimum of the curved mirror is shifted from the center of the mirror.
Since we have no ability to control mirror pitch angle once it is glued on the prism, the location of the curvature minima
should be characterized so that we can oush all of the misalignment in the horizontal direction.

Measurement technique:

When a curved mirror is completely axisymmetric (in terms of the mirror shape), any rotation of the mirror does not induce change on the axis of the refected beam.
If the curvature minimum is deviated from the center of the mirror, the reflected beam suffer precession. As we want to precisely rotate the mirror, we use the gluing
fixture for the PZT assembly. In this method, the back surface of the curved mirror is pushed on the mounting prism, and the lateral position of the mirror is precisely
defined by the fixture. As you rotate the mirror in clockwise viewing from the front, the spot moves in counter clockwise on the CCD.
curved_mirror_precession1.png

 

 

Setup and procedure:

The mounting prism (#21) is placed on the gluing fixture. A curved mirror under the test is loaded in the fixture with no PZT.
i.e. the back surface is aligned by the mounting prism. The fixing pressure is applied to the curved mirror by the front plate
with spring loads. The mirror needs be pushed from the top at least once to keep its defined position in the fixture.
The incident beam is slightly slated for the detection of the reflected spot. The beam is aligned and hits the center of the mirror as much as possible.

curved_mirror_precession2.png

The position of the spot on the CCD (WinCamD) is recorded, while the mirror is rotated 90deg at once. The rotation of the mirror is defined as shown in the figure below.
The angle origin is defined by the arrow mark of the mirror and rotated in clockwise being viewed from the front face. The mirror is rotated 540deg (8points) to check
the reproducibility.

curved_mirror_precession3.png

Measurement result:

8 point for each mirror is fitted by a circle. The fitting result provides the origin and radius of the circle, and the angle correspond to mirror angle of 0deg.

Analysis:

d: distance of the curvature minimum and the mirror center (quantity to be delived)

D: distance of the prove beam spot from the center of the mirror

R: Radius of curvature of the mirror

theta_R: angle of incidence/reflection

curved_mirror_precession4.png

 

 

The interesting consequence is that precession diameter (X-X') on the CCD does not depend on the spot position on the mirror.
This ensures the precision of the measurement. In the measurement, the radius of the precession (r = (X-X')/2) is obtained.

Therefore,

d = r R / (2 L)

Mirror name, distance[mm]
C1: 0.95
C3: 1.07
C4: 1.13
C5: 0.97
C6: 0.73
C7: 1.67
C8: 2.72
C9: 1.05
C10: 0.41
C11: 0.64
C12: 0.92
C13: 0.14

Resolution:
The angle to be rotated is depicted in the following plot for each mirror.

curved_mirror_minimum_position.png

Attachment 5: curved_mirror_precession.pdf
curved_mirror_precession.pdf
  90   Mon Apr 1 10:28:03 2013 KojiMechanicsGeneralAdditional UV blast for the top surface

[Koji, Lisa, Jeff, Zach]

Jeffs concern after talking with the glue company (EMI) was that the UV blast for the top side was not enough.

First we wanted to confirm if too much blasting is any harmful for the glue joint.

We took a test joint of FS-FS with the UV epoxy. We blasted the UV for 1min with ~15mm distance from the joint.
After the observation of the joint, we continued to blast more.
In total, we gave additional 5min exposure. No obvious change was found on the joint.

P3283483.JPG

Then proceed to blast the OMC top again. We gave 1 min additional blast on each glue joint.

 P3283459.jpg P3283473.JPG

  89   Mon Apr 1 03:23:48 2013 KojiOpticsGeneralUV power calibration

[Koji Lisa Jeff Zach]

Eric G bought a UV power meter from American Ultraviolet.

Our UV illuminator was calibrated by this power meter.

The first blast (i.e. cold start): 3.9W/cm^2

After many blasting: 8.3W/cm^2

The spec is 20W/cm^2

  88   Mon Apr 1 03:13:41 2013 KojiOpticsGeneralFailure of PZT-glass joints

[Koji, Jeff, Zach, Lisa]

We glued a test PZT-mirror assembly last week in order to make sure the heat cure of the epoxy does not make any problem
on the glass-PZT joints. The assembly was sent to Bob for the heat treatment. We received the assembly back from Bob on Wednesday.

We noticed that the assembly after the heat cure at 100degC had some voids in the epoxy layer
(looking like the fused silica surface was only 70% "wetted" by the epoxy).
The comparison of the assembly before and after the heat treatment is found in the slideshow at the bottom of the entry.

Initially our main concern was the impact to the control and noise performance.
An unexpected series resonance on the PZT transfer function and unwanted noise creation by the imperfect bonding may terribly ruin the IFO sensitivity.
In reality, after repeated poking by fingers, the PZT-prism joint was detached. This isn't good at all.
Note that there is no sign of degradation on the glass-glass joint.

We investigated the cause of this like:
- Difference of thermal expansion (3ppm/C PZT vs 0.55ppm/C fused silica)
- Insufficient curing of epoxy by UV (but this is the motivation of the heat cure)

Our resolution up to this point is to switch the glue to EP30-2. This means we will go through the heat cure test again.
Unfortunately there is no EP30-2 in stock at Caltech. We asked MIT to send us some packets of EP30-2.

Hardness of the epoxies is another concern. Through the epoxy investigation, we learned from Noliac that the glue for the PZT
should not be too hard (stiff) so as not to constrain the deformation of the PZT. EP30-2 has Shore D Hardness of 75 or more,
while Optocast UV epoxy has 88, and EPOTEK Epoxies, which Noliac suggested for gluing, has ~65. This should also be
confirmed by some measurement.  We will also ask Master Bond if they have information regarding the effect of curing
temperature on the hardness of the epoxy.  EP30-2 can be cured anywhere between RT and 200F (it's service range is up to 300F).
However, the entire breadboard, with the curved mirror sub-assemblies, will need to be baked at 110C to cure the UV Bond epoxy. 
We hope that exposure to relatively higher temps doesn't harden the EP30-2. The EP30-2 data sheet recommends an epoxy
thickness of 80-120 microns which is much thicker than we would like.

We also don't have a way tocontrol the thickness; though we could add glass spheres to the epoxy to control the thickness.
The thickness of the EP30-2 used to bond the metal wire guide prism on the core optics is much thinner at 15-25 microns.

 

  87   Fri Mar 29 08:55:00 2013 ZachOpticsConfigurationBeam launched into fiber

Quote:

Lisa and I concurred that it felt like we had converged to the optimum alignment and polarization, which would mean that the lack of coupling is all from mode mismatch. Since the input mode is well collimated, it seems unlikely that we could be off enough to only get ~1% coupling. One possibility is that the collimator is not well attached to the fiber itself. Since the Rayleigh range within it is very small, any looseness here can be critical.

My hypothesis about the input-side collimator turned out to be correct.

I removed the fiber from the collimator and mount at the input side, and then injected the illuminator beam from this side. Since we already saw a nice (but dim) IR beam emerging from the output side the other night, it followed that that collimator was correctly attached. With the illuminator injected from the input side, I also saw a nice, collimated red beam emerging from the output. So, the input collimator was not properly attached during our previous attempts, leading to the abysmal coupling.

The problem is that the mount does not allow you to remove and reattach the fiber while the collimator is already attached, and the dimensions make it hard to fit your fingers in to tighten the fiber to the collimator once the collimator is in the mount. I disassembled the mount and found a way to attach/reattach the fiber that preserves the tight collimator contact. I will upload a how-to shortly.

With this fix, I was able to align the input beam and get decent coupling:

EOM path: ~70%

AOM path: ~50%

  86   Thu Mar 28 03:37:07 2013 ZachOpticsConfigurationTest setup input optics progress

[Lisa, Zach]

Last night (Tuesday), I finished setting up and aligning most of the input optics for the OMC characterization setup. See the diagram below, but the setup consists of:

  • Faraday isolator/polarization definition
  • HWP+PBS for power splitting into two paths:
    • EOM path
      • Resonant EOM for PDH sideband generation
      • Broadband EOM for frequency scanning
    • AOM path
      • Double-passed ~200-MHz Isomet AOM for subcarrier generation. NOTE: in this case, I have chosen the m = -1 diffraction order due to the space constraints on the table.
  • Recombination of paths on a 50/50 beam splitter---half of the power is lost through the unused port into a black glass dump
  • Coupler for launching dual-field beam into a fiber (to OMC)

OMC_test_setup_IO.jpg 2013-03-27_05.48.20.jpg

Today, we placed some lenses into the setup, in two places:

  1. In the roundabout section of the AOM path that leads to the recombination, to re-match the AOM-path beam to that of the EOM path
  2. After the recombination beam splitter, to match the combined beam mode into the fiber

We (Koji, Lisa, and myself) had significant trouble getting more than ~0.1% coupling through the fiber, and after a while we decided to go to the 40m to get the red-light fiber illuminator to help with the alignment.

Using the illuminator, we realigned the input to the coupler and eventually got much better---but still bad---coupling of ~1.2% (0.12 mW out / 10 mW in). Due to the multi-mode nature of the illuminator beam, the output cannot be used to judge the collimation of the IR beam; it can only be used to verify the alignment of the beam.

With 0.12 mW emerging from the other end of the fiber, we could see the output quite clearly on a card (see photo below). This can tell us about the required input mode. From the looks of it, our beam is actually focused too strongly. We should probably replace the 75mm lens again with a slightly longer one.

Lisa and I concurred that it felt like we had converged to the optimum alignment and polarization, which would mean that the lack of coupling is all from mode mismatch. Since the input mode is well collimated, it seems unlikely that we could be off enough to only get ~1% coupling. One possibility is that the collimator is not well attached to the fiber itself. Since the Rayleigh range within it is very small, any looseness here can be critical.

2013-03-28_02.29.36.jpg

I think there are several people around here who have worked pretty extensively with fibers. So, I propose that we ask them to take a look at what we have done and see if we're doing something totally wrong. There is no reason to reinvent the wheel.

  85   Wed Mar 27 20:55:10 2013 KojiGeneralGeneralOMC building plan / procedure ~ Mar 28 Thu

28 March (Thu):

  • Rebuild the bottom template in the lab
  • Place the bottom template on the OMC
  • Glue the PZTs on the mounting prisms (x2)
  • Glue the curved mirrors on the PZTs (x2)
  • R&T measurement
  • Placing optics on the OMC breadboard
     
  • Better coupling to the fiber
  • Matching to the OMC cavity

 

29 March (Fri):

Start gluing bottom side: Set 4 cavity mirrors and 1 HR mirror
   and try to resonate beam.  Glue when OK.


Place BS and DCPD mounting brackets.  Glue when OK.
Friday: Place QPDs and rest of optics.  Glue when OK.

WB 1 April

  • Testing at CIT
    • Transmission / Coupling / Loss
    • FSR / TMS
    • Power dependence
    • PZT position dependence
    • Back scattering
    • Openloop TF
    • PZT TF
    • Noise measurement
  • Epoxy cure bake at CIT
  • Retest at CIT
    • ditto

WB 8 April or after

  • Packing
  • Shipping - Shipping box?
  • Optical Testing at LLO (2 days anticipated)
  84   Wed Mar 27 20:54:54 2013 KojiGeneralGeneralOMC building plan / procedure ~ Mar 27 Wed

27 March (Wed)

- AOM drift investigation (Lisa, Zach)

- Cavity input optics ~ Fiber coupling (Zach)

Action Items

  • Glue curved mirror sub-assys.
  • R&T measurement
  83   Wed Mar 27 20:54:45 2013 KojiGeneralGeneralOMC building plan / procedure ~ Mar 26 Mon

26 March (Tue):

- Curved mirror characterization (Koji, done)

- Input optics for the cavity locking (Zach)

Faraday, BB EOM, Resonant EOM, AOM, MZ

  82   Tue Mar 26 22:33:07 2013 KojiGeneralGeneralLoan for the OMC building

Loan from PSL Lab

- 300mm mirror with Ultima mount and pedestal
- Isopropanol small glass bottleReturned on Apr 12 2013.
- Newport 422-1S single-axis stage
Returned on Apr 12 2013.

Loan from ATF Lab

- 50/50 Cube BS 05BC16NP.9 without mount
- 1.5" pedestal (1/4-20 thread), 1/4" shim (1/4-20 through-hole), 1/8" shim (1/4-20 through-hole): 2 eachReturned on Aug 22, 2013.
- PBS & PBS mount
- Newport 422-1S single-axis stageReturned on Apr 12 2013.
- 10 ft BNC cable x 2
- some more BNC (labeled as ATF)Returned on May 20 2013.
- Y1-1037-45P with ultima mount 3inch post
- 1x Newfocus 5104 mirrorReturned on Aug 22, 2013.
- 4x ForkReturned on Aug 22, 2013.

Loan from 40m

- 4 BNC Ts and 1 BNC Ys
- 4 BNC Ts and 1 BNC YsReturned on May 20 2013.
- 6 BNC cablesReturned on May 20 2013.

- SONY CCD / CCD Monitor / CCD power supply
- Optical fiber tester (for fiber alignment) Returned on Apr 9 2013.

  81   Mon Mar 25 19:31:16 2013 KojiOpticsGeneralOMC Top-side gluing

[Koji Jeff Zach]

AAA

P3253372.jpg

BBB

P3253353.jpg

CCC

P3253394.jpg

DDD

P3253400.jpg

  80   Mon Mar 25 18:34:25 2013 KojiGeneralGeneralOMC building plan / procedure ~ Mar 25 Mon

25 March (Mon):

Inspect the test PZT assembly

  • => Give it to Bob. (done)

Glue topside components (done)

  • Clean up the table for the gluing work.
  • Prepare the transport fixture on the table.
  • Glass breadboard
    • Pick breadboard #1 (cf. [ELOG 27])
    • Wipe the entire glass breadboard with IPA
    • Place the breadboard in the fixture (check which is the upper side)
  • Gluing
    • Set the gluing template on the breadboard.
    • Place all of the glass components on the plate (just for confirmation)
    • Wipe (locally) both surfaces to be glued.
    • Apply glue on the component to be glued
    • Align the components in the template. Use the cantilever pusher if necessary.
    • Illuminate UV
    • Repeat the above process for all of the components.
  • Close the transport fixture and wrap with Al foil
  79   Mon Mar 25 02:04:05 2013 KojiGeneralGeneralOMC building plan / procedure ~ WB18

WB 18 March

  • Diode test
    • Dark current / Dark noise / Impedance
    • Quantum Efficiency test (but with glass)
    • Diode given to Bob for cleaning
  • Research possible issue of UV light on 1064 HR coating
    • ~ppm order loss increase after depositing 3J/cm^2 in 8 hours (i.e. same order to our illumination but in 10s for us)
    • Sent an e-mail to Ke-Xun Sun -> So far, no reply.
  • Glue test of PZT+prism+curved mirror with UV bond epoxy 
    • Done. Found some handling issues on the fixture.
  • Preparation of N2 line:
    • Done

Action items:

  • Bake test at 100°C for 1 hour at CIT 
    • Will be done on 25 Mon-26 Tue at Bob's lab
  • Curved mirror characterization
  • R&T measurement
  78   Sat Mar 23 16:36:15 2013 KojiElectronicsCharacterizationDiode QE measurement

Quantum efficiencies of the C30665GH diodes were measured. 

- The diode was biased by the FEMTO preamplifier.

- Diode Pin 1 Signal, Pin 2 +5V, Pin 3 open

- Preamp gain 10^3 V/A

- Beam power was measured by the thorlabs power meter.

 

PD #1
Incident: 12.82 +/- 0.02 mW
Vout: 9.161 +/- 0.0005 V
PD Reflection (Prompt): 0.404 mW
PD Reflection (Total): 1.168 mW

PD #2
Incident: 12.73 +/- 0.02 mW
Vout: 9.457 +/- 0.0005 V
PD Reflection (Prompt): 0.364 mW
PD Reflection (Total): 0.937 mW

PD #3
Incident: 12.67 +/- 0.02 mW
Vout: 9.1139 +/- 0.01 V
PD Reflection (Prompt): 0.383 mW
PD Reflection (Total): 1.272 mW

PD #4
Incident: 12.71 +/- 0.02 mW
Vout: 9.3065 +/- 0.0005 V
PD Reflection (Prompt): 0.393 mW
PD Reflection (Total): 1.033 mW

PD #5
Incident: 12.69 +/- 0.02 mW
Vout: 9.1071 +/- 0.005 V
PD Reflection (Prompt): 0.401 mW
PD Reflection (Total): 1.183 mW

PD #6
Incident: 12.65 +/- 0.02 mW
Vout: 9.0310 +/- 0.01 V
PD Reflection (Prompt): 0.395 mW
PD Reflection (Total): 1.306 mW

PD #7
Incident: 12.67 +/- 0.02 mW
Vout: 9.0590 +/- 0.0005 V
PD Reflection (Prompt): 0.411 mW
PD Reflection (Total): 1.376 mW

PD #8
Incident: 12.63 +/- 0.01 mW
Vout: 9.0790 +/- 0.0005 V
PD Reflection (Prompt): 0.420 mW
PD Reflection (Total): 1.295 mW

PD #9
Incident: 12.67 +/- 0.02 mW
Vout: 9.2075 +/- 0.0005 V
PD Reflection (Prompt): 0.384 mW
PD Reflection (Total): 1.091 mW

PD #10
Incident: 12.70 +/- 0.01 mW
Vout: 9.0880 +/- 0.001 V
PD Reflection (Prompt): 0.414 mW
PD Reflection (Total): 1.304 mW

PD #11
Incident: 12.64 +/- 0.01 mW
Vout: 9.2861 +/- 0.0005 V
PD Reflection (Prompt): 0.416 mW
PD Reflection (Total): 1.152 mW

PD #12
Incident: 12.68 +/- 0.02 mW
Vout: 9.3650 +/- 0.001 V
PD Reflection (Prompt): 0.419 mW
PD Reflection (Total): 1.057 mW

PD #13
Incident: 12.89 +/- 0.01 mW
Vout: 9.3861 +/- 0.001 V
PD Reflection (Prompt): 0.410 mW
PD Reflection (Total): 1.047 mW

 

PD serial number
 1 - 0782
 2 - 1139
 3 - 0793
 4 - 0732
 5 - 0791
 6 - 0792
 7 - 0787
 8 - 0790
 9 - 0781
10 - 0784
11 - 1213
12 - 1208
13 - 1209

 

{
  {1, 12.82, 9.161, 0.404, 1.168},
  {2, 12.73 , 9.457, 0.364 , 0.937} ,
  {3, 12.67 , 9.1139, 0.383 , 1.272 },
  {4, 12.71 , 9.3065, 0.393 , 1.033 },
  {5, 12.69, 9.1071, 0.401 , 1.183 },
  {6, 12.65, 9.0310, 0.395 , 1.306} ,
  {7, 12.67, 9.0590, 0.411 , 1.376} ,
  {8, 12.63 , 9.0790, 0.420 , 1.295} ,
  {9, 12.67 , 9.2075, 0.384 , 1.091} ,
  {10, 12.70, 9.0880, 0.414 , 1.304 },
  {11, 12.64 , 9.2861, 0.416 , 1.152} ,
  {12, 12.68 , 9.3650, 0.419 , 1.057} ,
  {13, 12.89 , 9.3861, 0.410 , 1.047}
};

Attachment 1: P3213308.JPG
P3213308.JPG
Attachment 2: P3213310.JPG
P3213310.JPG
  77   Sat Mar 23 13:34:14 2013 KojiOpticsGeneralPZT assembly prototype glued

Prototype PZT assembly

Motivation:

Before we glue the PZT assembly, we need to build a prototype. This is to confirm the heat cure process
does not cause any cracking of the PZT or glass components. The CTE of the PZT is 2~3ppm
(depends on the direction) while the one for Fused Silica is 0.55ppm.

Materials:

- A fused silica substrate, 1/2" in dia. Supplied from Garilynn. I defined the chamfered side as the front side.

- PZT: Noliac NAC2124, serial #24, this is a spare PZT as this has the worst length to angle coupling.

- Mounting Prism: D1102069 SN22. This has the worst perpendicularity among the prisms.

- Fixtures:

D1300185 aLIGO OMC CURVED MIRROR BONDING FIXTURE ASSY
D1300186 aLIGO OMC CURVED MIRROR BONDING FIXTURE FRONT
D1300187 aLIGO OMC CURVED MIRROR BONDING FIXTURE BACK
D1300188 aLIGO OMC CURVED MIRROR BONDING FIXTURE RING

P3223322.jpg

Procedure:

- Wipe all of the components with the isopropanol.

- Attach the back piece of the fixture on the Al wrapped bracket.
(The current 4-40 screws for the middle piece are too long and stick out from the back side of the back piece.
Therefore a 1/16" shim for a 1/2" rod is inserted between the bracket and the back piece)

- Brought a glue package to the lab (10:40PM)

- Loosely attach the middle piece to the back piece with four 4-40 screws.

- Insert the mounting prism in the fixture. Insert the PZT in the fixture too.

- Insert a dummy substrate in the fixture.

- Attach the front piece with spring loaded screws.

- Align the PZT and the optic in the fixture. (Basically apply downward force to them)

- Test the rigidity of the assembly (11:30PM)

- Remove the PZT and the mirror. Apply UV epoxy.
(A single dub was applied for each PZT surface of the PZT but this was too much.)

- Make sure the PZT and the optic are aligned by applying the downward force.

- Illuminate UV light from the front.

- Illuminate UV light from the back. (11:50PM)

Procedural issues:

- Long 4-40 screws (described above)
(Circumvented)

- As the PZT is not constrained with the middle piece, it tends to move vertically and rotationally
because of the wire tension. (This is not a mistake but the design so that the PZT is constrained by the optic.)
Therefore after applying glue on the PZT, the motion of the PZT spreads the glue on the back surface of
the curved mirror.

(Solution to be tried) Our solution is to glue the PZT and the mounting prism first with a dummy optics (made of SF2).
The wires should be tacked somewhere on the mount 

- The amount of glue on the PZT was too much. I gave one dub of glue for each side.
As a result, excess glue leaks out along the ring.

- The front plate has a chamfered hole but this tends to slip and move the mirror vertically.
Later I used the flat side of the plate to hole the mirror.
(Circumvented) It seems that this hold the mirror in a better way as the plate can't rock

- Spring load for the front plate was too strong. This was because the natural length of the spring was too long.
(Circumvented) The spring was cut at the length of the 4-40 screw. Then attaching the screws became completely fine.

P3223323.jpg


Result:

P3233336.jpg P3233348.jpg

Slide show:

  76   Sat Mar 23 02:41:00 2013 KojiOpticsGeneralBlack glass beam dumps for the first OMC

Received black glass beam dumps from MIT

- gluing by EP30-2 looks pretty fine. Enough sturdy.

- some gap visible between the glass => incident angle should be considered so that the first beam does not exit from the gap

- Dusts are visible on the glass surface. Some have a lot, the other have less. But every piece still needs to be wiped.

Attachment 1: P3223314.JPG
P3223314.JPG
Attachment 2: P3223319.jpg
P3223319.jpg
Attachment 3: P3223320.jpg
P3223320.jpg
  75   Sat Mar 23 02:32:23 2013 KojiFacilityGeneralN2 cylinder delivered

Preparation for ionized N2 blow

- 99.9998% N2 cylinder delivered (ALPHAGAZ 2 grade by AIR LIQUIDE) ALPHAGAZ 2 [PDF]

- Filter and Arcing module already in the lab

- A brass regulator to be installed (Done - March 24)

- 50 ft air line already in the lab / needs to be wiped/rinsed (Done - March 24)

- Air line and filter installed (Done - March 24)

Attachment 1: P3233349.jpg
P3233349.jpg
  74   Wed Mar 20 09:38:02 2013 ZachOpticsCharacterization[LLO] OMC test bench modified

 For various reasons, I had to switch NPROs (from the LightWave 126 to the Innolight Prometheus).

I installed the laser, realigned the polarization and modulation optics, and then began launching the beam into the fiber, though I have not coupled any light yet.

A diagram is below. Since I do not yet have the AOM, I have shown that future path with a dotted line. Since we will not need to make AMTFs and have a subcarrier at the same time, I have chosen to overload the function of the PBS using the HWP after the AEOM. We will operate in one of two modes:

  1. AMTF mode: The AOM path is used as a beam dump for the amplitude modulation setup. A razor dump should physically be placed somewhere in the AOM path.
  2. Subcarrier mode: The AEOM is turned off and the HWP after it is used to adjust the carrier/subcarrier power ratio. I chose a 70T / 30R beamsplitter for the recombining, since we want to be able to provide ~100 mW with the carrier for transmission testing, and we don't need a particularly strong subcarrier beam for probing.

new_setup_traced.jpg

One thing that concerns me slightly: the Prometheus is a dual-output (1064nm/532nm) laser, with separate ports for each. I have blocked and locked out the green path physically, but there is some residual green light visible in the IR output. Since we are planning to do the OMC transmission testing with a Si-based Thorlabs power meter---which is more sensitive to green than IR---I am somewhat worried about the ensuing systematics. I *think* we can minimize the effect by detuning the doubling crystal temperature, but this remains to be verified.

 EDIT (ZK): Valera says there should be a dichroic beam splitter in the lab that I can borrow. This should be enough to selectively suppress the green.

  73   Sun Mar 17 21:59:47 2013 KojiElectronicsCharacterizationDiode testing ~ DCPD

- For the dark noise measurement, the lid of the die-cast case should also contact to the box for better shielding. This made the 60Hz lines almost completely removed, although unknown 1kHz harmonics remains.

- The precise impedance of the setup can not be obtained from the measurement box; the cable in between is too long. The diode impedance should be measured with the impedance measurement kit.

- With the impedance measurement kit, the bias voltage of +5V should be used, in stead of -5V.

- diode characteristics measured at 10-100MHz

- Typical impedance characteristics of the diodes

Excelitas (Perkin-Elmer) C30665GH Rs=9Ohm, Cd=220pF, L=0~1nH (Vr=5V)

Excelitas (Perkin-Elmer) C30642G Rs=12Ohm, Cd=100pF, L=~5nH (Vr=5V) longer thin wire in a can?

Excelitas (Perkin-Elmer) C30641GH Rs=8Ohm, Cd=26pF, L=12nH (Vr=5V) leg inductance? (leg ~30mm)

- PD serial

C30665GH, Ls ~ 1nH

  1 - 0782 from PK, Rs=8.3Ohm, Cd=219.9pF
  2 - 1139 from PK, Rs=9.9Ohm, Cd=214.3pF
  3 - 0793 from PK, Rs=8.5Ohm, Cd=212.8pF
  4 - 0732 from PK, Rs=7.4Ohm, Cd=214.1pF
  5 - 0791 from PK, Rs=8.4Ohm, Cd=209.9pF
  6 - 0792 from PK, Rs=8.0Ohm, Cd=219.0pF
  7 - 0787 from PK, Rs=9.0Ohm, Cd=197.1pF
  8 - 0790 from PK, Rs=8.4Ohm, Cd=213.1pF
  9 - 0781 from PK, Rs=8.2Ohm, Cd=216.9pF
10 - 0784 from PK, Rs=8.2Ohm, Cd=220.0pF
11 - 1213 from the 40m, Rs=10.0Ohm, Cd=212.9pF
12 - 1208 from the 40m, Rs=9.9Ohm, Cd=216.8pF
13 - 1209 from the 40m, Rs=10.0Ohm, Cd=217.5pF

C30642G, Ls ~ 12nH

20 - 2484 from the 40m EG&G, Rs=12.0Ohm, Cd=99.1pF
21 - 2487 from the 40m EG&G, Rs=14.2Ohm, Cd=109.1pF
22 - 2475 from the 40m EG&G glass crack, Rs=13.5Ohm, Cd=91.6pF
23 - 6367 from the 40m ?, Rs=9.99Ohm, Cd=134.7pF
24 - 1559 from the 40m Perkin-Elmer GH, Rs=8.37Ohm, Cd=94.5pF
25 - 1564 from the 40m Perkin-Elmer GH, Rs=7.73Ohm, Cd=94.5pF
26 - 1565 from the 40m Perkin-Elmer GH, Rs=8.22Ohm, Cd=95.6pF
27 - 1566 from the 40m Perkin-Elmer GH, Rs=8.25Ohm, Cd=94.9pF
28 - 1568 from the 40m Perkin-Elmer GH, Rs=7.83Ohm, Cd=94.9pF
29 - 1575 from the 40m Perkin-Elmer GH, Rs=8.32Ohm, Cd=100.5pF

C30641GH, Perkin Elmer, Ls ~ 12nH

30 - 8983 from the 40m Perkin-Elmer, Rs=8.19Ohm, Cd=25.8pF
31 - 8984 from the 40m Perkin-Elmer, Rs=8.39Ohm, Cd=25.7pF
32 - 8985 from the 40m Perkin-Elmer, Rs=8.60Ohm, Cd=25.2pF
33 - 8996 from the 40m Perkin-Elmer, Rs=8.02Ohm, Cd=25.7pF
34 - 8997 from the 40m Perkin-Elmer, Rs=8.35Ohm, Cd=25.8pF
35 - 8998 from the 40m Perkin-Elmer, Rs=7.89Ohm, Cd=25.5pF
36 - 9000 from the 40m Perkin-Elmer, Rs=8.17Ohm, Cd=25.7pF

 

Note:
  1mm Au wire with dia. 10um -> 1nH, 0.3 Ohm
20mm BeCu wire with dia. 460um -> 18nH, 0.01 Ohm

Attachment 1: OMCPD_TEST_130317.pdf
OMCPD_TEST_130317.pdf OMCPD_TEST_130317.pdf OMCPD_TEST_130317.pdf OMCPD_TEST_130317.pdf OMCPD_TEST_130317.pdf OMCPD_TEST_130317.pdf OMCPD_TEST_130317.pdf OMCPD_TEST_130317.pdf
  72   Fri Mar 15 02:15:45 2013 KojiElectronicsCharacterizationDiode testing

Diode testing

o Purpose of the measurement

- Test Si QPDs (C30845EH) for ISC QPDs Qty 30 (i.e. 120 elements)

- Test InGaAs PDs (C30665GH) for OMC Qty 10 (i.e. 10 elements)

o Measurement Kit

- Inherited from Frank.

- Has relays in it.

- D0 and D1 switches the measurement instrument connected to an element

- D2 and D3 switches the element of the QPDs

- Digital switch summary

d0 d1 0 0 - ln preamp
d0 d1 1 0 - dark c
d0 d1 0 1 - omc preamp
d0 d1 1 1 - impedance

d2 d3 0 0 - A x x x
d2 d3 1 0 - C x o x
d2 d3 0 1 - B o x o
d2 d3 1 1 - D o o o

- The universal board in the box is currently configured for C30845.
  Pin1 - Elem A. Pin3 - B, Pin7 - C, Pin9 - D, Pin 12 - Case&Bias

o Labview interface

- Controls NI-USB-6009 USB DAQ interface and Agilent 82357B USB-GPIB interface

o Dark current measurement

- Borrowed Peter's source meter KEITHLEY 2635A

- For C30845GH the maxmum reverse bias is set to -20V. This drops the voltage of the each element to the bias voltage.

o Spectrum measurement

- The elements are connected to FEMTO LN current amp DLPCA-200.

- Bias voltage is set to +10V. This lifts up the outside of the amplifier input to +10V.

 

o Impedance measurement

- Agilent 4395A at PSL lab with impedance measurement kit

- For C30845GH the maxmum reverse bias is set to -15V. This drops the voltage of the each element to the bias voltage.

- Calibration: open - unplug the diode from the socket, short - use a piece of resister lead, 50Ohm - a thin metal resister 51Ohm

- Freq range: 30-50MHz where the response of the cables in the setup is mostly flat.

- Labview VI is configured to read the equivalent circuit parameters in the configuration "D" (series LCR).

- Labview fails to read the series resistance. This was solved by first read the equiv circuit param and then read it with Sim F-CHRST.
  F-CHRST does nothing on the parameters so the second request successfully acquires the first ones.

 

Attachment 1: QPD_GR_TEST_130316.pdf
QPD_GR_TEST_130316.pdf QPD_GR_TEST_130316.pdf QPD_GR_TEST_130316.pdf QPD_GR_TEST_130316.pdf QPD_GR_TEST_130316.pdf QPD_GR_TEST_130316.pdf QPD_GR_TEST_130316.pdf QPD_GR_TEST_130316.pdf
  71   Thu Mar 14 22:18:23 2013 KojiGeneralGeneralNew loans for the diode test

ALL returned

Loan from ATF:

2 blue banana cables returned on Jun 4, 2013

BNC cable returned on Mar 21, 2013

TENMA triple power supply returned on July 17, 2015

From 40m:

4x GPIB cables returned on Mar 21, 2013

From EE shop:

red banana cables returned on Jun 4, 2013

  70   Thu Mar 14 17:06:21 2013 KojiMechanicsGeneralOMC SUS work @LLO

EDIT (ZK): All photos on Picasa. Also, I discovered that since Picasa was migrated to Google+ only,
you no longer have the option to embed a slideshow like you used to. Lame, Google.

Photos sent from Zach

(3D VIEW)

2013-03-14_16.04.07.jpg2013-03-14_16.03.40.jpg

  69   Thu Mar 7 15:53:47 2013 KojiMechanicsGeneralOMC Transportation fixture, OMC PD/QPD mounts

P3073218.JPG 

P3073219.JPG

P3073223.JPG

P3073225.JPG

P3073227.JPG

  68   Wed Mar 6 23:24:58 2013 ZachOpticsCharacterizationeLIGO OMC visibility vs. power measurement details

I found that, in fact, I had lowered the modulation depth since when I measured it to be 0.45 rads --> Psb = 0.1.

Here is the sweep measurement:

TEK00005.PNG TEK00007.PNG

This is Psb = 0.06 --> gamma = 0.35 rads.

This changes the "raw transmission" and "coupling", but not the inferred visibility:

eOMC_visibility_3_4_13.png

I also measured the cavity AMTF at three powers today: 0.5 mW, 10 mW, and 45 mW input.

eOMC_AMTF_vs_power_3_6_13.png

They look about the same. If anything, the cavity pole seems slightly lower with the higher power, which is counterintuitive. The expected shift is very small (~10%), since the decay rate is still totally dominated by the mirror transmissions even for the supposed high-loss state (Sam and Sheon estimated the roundtrip loss at high power to be ~1400 ppm, while the combined coupling mirrors' T is 1.6%). I have not been able to fit the cavity poles consistently to within this kind of error.

  67   Tue Mar 5 19:37:00 2013 ZachOpticsCharacterizationeLIGO OMC visibility vs. power measurement details

EDIT (ZK): Koji points out that (1 - Ti) should really be the non-resonant reflectivity of the aligned cavity, which is much closer to 1. However, it should *actually* be the non-resonant reflectivity of the entire OMC assembly, including the steering mirror (see bottom of post). The steering mirror has T ~ 0.3%, so the true results are somewhere between my numbers and those with (1 - Ti) -> 1. In practice, though, these effects are swamped by the other errors.

More information about the power-dependent visibility measurement:

As a blanket statement, this measurement was done by exact analogy to those made by Sam and Sheon during S6 (c.f. LHO iLog 11/7/2011 and technical note T1100562), since it was supposed to be a verification that this effect still remains. There are absolutely better ways to do (i.e., ways that should give lower measurement error), and these should be investigated for our characterization. Obviously, I volunteer.

All measurements were made by reading the output voltages produced by photodetectors at the REFL and TRANS ports. The REFL PD is a BBPD (DC output), and the TRANS is a PDA255. Both these PDs were calibrated using a Thorlabs power meter (Controller: PM100D; Head: S12XC series photodiode-based---not sure if X = 0,2... Si or Ge) at the lowest and highest power settings, and these results agreed to the few-percent level. This can be a major source of error.

The power was adjusted using the HWP/PBS combination towards the beginning of the experiment. For reference, an early layout of the test setup can be seen in LLO:5978 (though, as mentioned above, the REFL and TRANS PDs have been replaced since then---see LLO:5994). This may or may not be a "clean" way to change the power, but the analysis should take the effect of junk light into account.

eOMC_visibility_3_4_13.png

Below is an explanation of the three traces in the plot. First:

  • TRANS: TRANS signal calibrated to W
  • REFL_UL: REFL signal while cavity is unlocked, calibrated to W
  • REFL_L: REFL signal while cavity is locked, calibrated to W
  • Psb: Sideband power (relative to carrier)
  • Ti: Input mirror transmission (in power)

Now, the traces

  1. Raw transmission: This measurement is simple. It is just the raw throughput of the cavity, corrected for the power in the sidebands which should not get through. I had the "AM_REF" PD, which could serve as an input power monitor, but I thought it was better to just use REFL_UL as the input power monitor and not introduce the error of another PD. This means I must also correct for the reduction in the apparent input power as measured at the REFL PD due to the finite transmission of the input coupler. This was not reported by Sam and Sheon, but can be directly inferred from their data.
    • trans_raw = TRANS ./ ( REFL_UL * (1 - Psb) * (1 - Ti) )
    • Equivalently, trans_raw = (transmitted power) ./ (input power in carrier mode)
  2. Coupling: This is how much of the power incident on the cavity gets coupled into the cavity (whether it ends up in transmission or at a loss port). Sheon plots something like (1 - coupling) in his reply to the above-linked iLog post on 11/8/2011.
    • coupling = ( REFL_UL * (1 - Ti) - REFL_L ) ./ ( REFL_UL * (1 - Psb) * (1 - Ti) )
    • Equivalently, coupling = [ (total input power) - (total reflected power on resonance) ] ./ (input power in carrier mode)
  3. Visibility: How much of the light that is coupled into the cavity is emerging from the transmitted port? This is what Sam and Sheon call "throughput" or "transmission" and is what is reported in the majority of their plots.
    • visibility = TRANS ./ ( REFL_UL * (1 - Ti) - REFL_L )
    • Equivalently, visibility = (transmitted power) ./ [ (total input power) - (total reflected power on resonance) ]
    • Also equivalently, visibility = trans_raw ./ coupling

The error bars in the measurement were dominated, roughly equally, by 1) systematic error from calibration of the PDs with the power meter, and 2) error from noise in the REFL_L measurement (since the absolute AC noise level in TRANS and REFL_L is the same, and TRANS >> REFL_L, the SNR of the latter is worse).

(1) can be helped by making ALL measurements with a single device. I recommend using something precise and portable like the power meter to make measurements at all the necessary ports. For REFL_L/UL, we can place a beam splitter before the REFL PD, and---after calibrating for the T of this splitter very well using the same power meter---both states can be measured at this port.

(2) can probably be helped by taking longer averaging, though at some point we run into the stability of the power setting itself. Something like 30-60s should be enough to remove the effects of the REFL_L noise, which is concentrated in the few-Hz region in the LLO setup.

One more thing I forgot was the finite transmission of the steering mirror at the OMC input (the transmission of this mirror goes to the QPDs). This will add a fixed error of 0.3%, and I will take it into account in the future.

  66   Fri Mar 1 23:52:18 2013 KojiOpticsCharacterizationWedge measurement with the autocollimator and the rotation stage

Measurement:

  • E1:   α = 0.672 deg, β = +0.0 arcmin (0 div up)
  • E2:   α = 0.631 deg, β = - 0.3 arcmin (-0.15 div down)
  • E3:   α = 0.642 deg, β = +0.0 arcmin (0 div up)
  • E4:   α = 0.659 deg, β = +1.4 arcmin (0.7 div up)
  • E5:   α = 0.695 deg, β = +0.5 arcmin (0.5 div up)
  • E6:   α = 0.665 deg, β = - 0.4 arcmin (-0.2 div down)
  • E7:   α = 0.652 deg, β = +1.0 arcmin (0.5 div up)
  • E8:   α = 0.675 deg, β = +2.0 arcmin (1.0 div up)
  • E9:   α = 0.645 deg, β = - 2.4 arcmin (-1.2 div down)
  • E10: α = 0.640 deg, β = +2.2 arcmin (1.1 div up)
  • E11: α = 0.638 deg, β = +1.6 arcmin (0.8 div up)
  • E12: α = 0.660 deg, β = +1.6 arcmin (0.8 div up)
  • E13: α = 0.638 deg, β = +0.8 arcmin (0.4 div up)
  • E14: α = 0.655 deg, β = +0.4 arcmin (0.2 div up)
  • E15: α = 0.640 deg, β = +1.4 arcmin (0.7 div up)
  • E16: α = 0.655 deg, β = +0.6 arcmin (0.3 div up)
  • E17: α = 0.650 deg, β = +0.8 arcmin (0.4 div up)
  • E18: α = 0.640 deg, β = +2.4 arcmin (1.2 div up)

Analysis:

  • \theta_H = ArcSin[Sin(α) / n]
  • \theta_V = ArcSin[Sin(β) / n]/2
     
  • E1:   \theta_H = 0.460 deg, \theta_V =   0.000 deg
  • E2:   \theta_H = 0.432 deg, \theta_V =  -0.0034 deg
  • E3:   \theta_H = 0.439 deg, \theta_V =   0.000 deg
  • E4:   \theta_H = 0.451 deg, \theta_V =  0.016 deg
  • E5:   \theta_H = 0.475 deg, \theta_V =  0.011 deg
  • E6:   \theta_H = 0.455 deg, \theta_V =  -0.0046 deg
  • E7:   \theta_H = 0.446 deg, \theta_V =  0.011 deg
  • E8:   \theta_H = 0.462 deg, \theta_V =  0.023 deg
  • E9:   \theta_H = 0.441 deg, \theta_V =  -0.027 deg
  • E10:   \theta_H = 0.438 deg, \theta_V = 0.025 deg
  • E11:   \theta_H = 0.436 deg, \theta_V = 0.018 deg
  • E12:   \theta_H = 0.451 deg, \theta_V = 0.018 deg
  • E13:   \theta_H = 0.436 deg, \theta_V = 0.0091 deg
  • E14:   \theta_H = 0.448 deg, \theta_V = 0.0046 deg
  • E15:   \theta_H = 0.438 deg, \theta_V = 0.016 deg
  • E16:   \theta_H = 0.448 deg, \theta_V = 0.0068 deg
  • E17:   \theta_H = 0.444 deg, \theta_V = 0.0091 deg
  • E18:   \theta_H = 0.438 deg, \theta_V = 0.027 deg
  65   Fri Mar 1 23:06:15 2013 KojiOpticsConfigurationMore perpendicularity test final

Perpendicularity of the "E" mirror was measured.


Mounting Prisms:
(criteria: 30arcsec = 145urad => 0.36mm spot shift)
SN  Meas.(div) ArcSec Spec.
10   0.3989    11.97   29    good
11   0.2202     6.60   16
    good
16   0.1907     5.72    5
    good
20  -0.591    -17.73    5
    good
21  -2.378    -71.34   15

21  -1.7      -51.     15
01  -0.5      -15.     52
02  -2.5      -75.     48
06  -1.0      -30.     15
    good
07   1.7       51.     59
12  -2.2      -66.     40
13  -0.3      - 9.     12
    good
14  -2.8      -84.     27
15  -2.5      -75.     50
17   0.7       21.     48
22   2.9       87.    
63

Mirror A:
A1  -0.5      -15.     NA    good
A3   0.5       15.     NA
    good
A4   0.9       27.     NA
    good
A5   0.4       12.     NA
    good
A6   0.1        3.
    NA    good
A7   0.0        0.
    NA    good
A8   0.0       
0.     NA    good
A9   0.0       
0.     NA    good
A10  1.0      
30.     NA    good
A11  0.3       
9.     NA    good
A12  0.1       
3.     NA    good
A13  0.0       
0.     NA    good
A14  0.6      
18.     NA    good

Mirror B:
B1  -0.9     
-27.     NA    good
B2  -0.6     
-18.     NA    good
B3  -0.9     
-27.     NA    good
B4   0.7      
21.     NA    good
B5  -1.1     
-33.     NA
B6  -0.6     
-18.     NA    good
B7  -1.8     
-54.     NA
B8  -1.1     
-33.     NA
B9   1.8      
54.     NA
B10  1.2       
36.     NA   
B11 -1.7     
-51.     NA
B12  1.1       
33.     NA

Mirror E:
E1  -0.8      -24.     NA    good
E2  -0.8      -24.    
NA    good
E3  -0.25     - 7.5   
NA    good
E4  -0.5      -15.     NA
    good
E5   0.8       24.     NA    good
E6  -1.0      -30.     NA
    good
E7  -0.2      - 6.     NA
    good
E8  -0.8      -24.     NA
    good
E9  -1.0      -30.     NA
    good
E10  0.0        0.     NA
    good
E11 -1.0      -30.     NA
    good
E12 -0.3      - 9.     NA
    good
E13 -0.8      -24.     NA
    good
E14 -1.0      -30.     NA    good
E15 -1.2      -36.     NA

E16 -0.7      -21.     NA
    good
E17 -0.8      -24.     NA
    good
E18 -1.0      -30.     NA
    good

  64   Wed Feb 27 18:18:48 2013 KojiOpticsConfigurationMore perpendicularity test

Mounting Prisms:
(criteria: 30arcsec = 145urad => 0.36mm spot shift)
SN  Meas.(div) ArcSec Spec.
10   0.3989    11.97   29    good
11   0.2202     6.60   16
    good
16   0.1907     5.72    5
    good
20  -0.591    -17.73    5
    good
21  -2.378    -71.34   15

21  -1.7      -51.     15
01  -0.5      -15.     52
02  -2.5      -75.     48
06  -1.0      -30.     15
    good
07   1.7       51.     59
12  -2.2      -66.     40
13  -0.3      - 9.     12
    good
14  -2.8      -84.     27
15  -2.5      -75.     50
17   0.7       21.     48
22   2.9       87.    
63

Mirror A:
A1  -0.5      -15.     NA    good
A3   0.5       15.     NA
    good
A4   0.9       27.     NA
    good
A5   0.4       12.     NA
    good
A6   0.1        3.
    NA    good
A7   0.0        0.
    NA    good
A8   0.0       
0.     NA    good
A9   0.0       
0.     NA    good
A10  1.0      
30.     NA    good
A11  0.3       
9.     NA    good
A12  0.1       
3.     NA    good
A13  0.0       
0.     NA    good
A14  0.6      
18.     NA    good

Mirror B:
B1  -0.9     
-27.     NA    good
B2  -0.6     
-18.     NA    good
B3  -0.9     
-27.     NA    good
B4   0.7      
21.     NA    good
B5  -1.1     
-33.     NA
B6  -0.6     
-18.     NA    good
B7  -1.8     
-54.     NA
B8  -1.1     
-33.     NA
B9   1.8      
54.     NA
B10  1.2       
36.     NA   
B11 -1.7     
-51.     NA
B12  1.1       
33.     NA

  63   Thu Feb 21 18:44:18 2013 KojiOpticsConfigurationPerpendicularity test

Perpendicularity test of the mounting prisms:

The perpendicularity of the prism pieces were measured with an autocollimator.

Two orthogonally jointed surfaces forms a part of a corner cube.
The deviation of the reflected image from retroreflection is the quantity measured by the device.

When the image is retroreflected, only one horizontal line is observed in the view.
If there is any deviation from the retroreflection, this horizontal line splits into two
as the upper and lower halves have the angled wavefront by 4x\theta. (see attached figure)

The actual reading of the autocollimator is half of the wavefront angle (as it assumes the optical lever).
Therefore the reading of the AC times 30 gives us the deviation from 90deg in the unit of arcsec.

SN / measured / spec

SN10: 12.0 arcsec (29 arcsec)

SN11: 6.6 arcsec (16 arcsec)

SN16: 5.7 arcsec (5 arcsec)

SN20: -17.7 arcsec (5 arcsec)

SN21: - 71.3 arcsec (15 arcsec)

 

Attachment 1: perpendicularity_test.pdf
perpendicularity_test.pdf perpendicularity_test.pdf
Attachment 2: P2203206.JPG
P2203206.JPG
  62   Thu Feb 7 23:01:45 2013 KojiOpticsCharacterizationUV epoxy gluing test

 [Jeff, Yuta, Koji]

Gluing test with UV-cure epoxy Optocast 3553-LV-UTF-HM

- This glue was bought in the end of October (~3.5 months ago).

- The glue was taken out from the freezer at 1:20pm.
- Al sheet was laid on the optical table. We made a boat with Al foil and pour the glue in it (@1:57pm)
- We brought two kinds of Cu wires from the 40m. The thicker one has the diameter of 1.62mm.
The thinner one has the diameter of 0.62mm. We decided to use thinner one being cut into 50mm in length.

- The OMC glass prisms have the footprint of 10mmx20mm = 200mm^2. We tested several combinations
of the substrates. Pairs of mirrors with 1/2" mm in dia. (127mm) and a pair of mirrors with 20mm in dia. (314mm).

- Firstly, a pair of 1/2" mirrors made of SF2 glass was used. A small dub on a thinner Cu wire was deposited on a mirror.
  We illuminated the glue for ~10sec. When the surfaces of the pair was matched, the glue did not spread on the entire
  surface. The glue was entirely spread once the pressure is applied by a finger. Glue was cured at 2:15pm. 12.873mm
  thickness after the gluing.

Some remark:
1. We should be careful not to shine the glue pot by the UV illuminator.
2. The gluing surface should be drag wiped to remove dusts on the surface.

- Secondly, we moved onto 20mm mirror pair taken from the remnant of the previous gluing test by the eLIGO people.
This time about 1.5 times more glue was applied.

- The third trial is to insert small piece of alminum foil to form a wedge. The thickness of the foil is 0.041mm.
The glue was applied to the pair of SF2 mirror (1/2" in dia.). A small dub (~1mm in dia) of the glue was applied.
The glue filled the wedge without any bubble although the glue tried to slide out the foil piece from the wedge.
So the handling was a bit difficult. After the gluing we measured the thickness of the wedge by a micrometer gauge.
The skinny side was 12.837mm, and the thicker side was 12.885mm. This is to be compared with the total thickness
12.823mm before the gluing. The wedge angle is 3.8mrad (0.22deg). The glue dub was applied at 2:43, and the UV
illumination was applied at 2:46.

- At the end we glued a pair of fused silica mirrors. The total thickness before the gluing was 12.658 mm.
The glue was applied at 2:59pm. The thickness after the gluing is 12.663 mm.
This indicates the glue thickess is 5um.

  61   Thu Feb 7 21:35:46 2013 KojiGeneralGeneralDmass's loan of LB1005 / A2&C2 sent to Fullerton / First Contact @40m

Dmass borrowed the LB1005 servo amp from the OMC lab.
It happened this week although it seems still January in his head.
Got it back on Mar 24th

The A2 and C2 mirrors have been sent to Josh Smith at Fullerton for the scatterometer measurement.

First Contact kit (incl. Peek Sheets)
Manasa borrowed the kit on Feb 7.
Got it back to the lab.

Attachment 1: P2053119.jpg
P2053119.jpg
Attachment 2: P2053120.JPG
P2053120.JPG
ELOG V3.1.3-